OpenStudy (anonymous):

H = 65 - m solve for m

OpenStudy (anonymous):

65-H

OpenStudy (saifoo.khan):

m = 65 - H

OpenStudy (anonymous):

The answer is in the back of the book...how do I go about solving..anyone know the steps?

jimthompson5910 (jim_thompson5910):

H = 65 - m ... Start with the given equation H+m = 65 - m+m ... Add m to both sides H+m = 65 ... Combine like terms H+m-H = 65-H ... Subtract H from both sides m = 65-H ... Combine like terms.

OpenStudy (anonymous):

Thank you! I have one more.. p=2/mtq solve for q

jimthompson5910 (jim_thompson5910):

Is it \[\large p=\frac{2}{mtq}\] ???

OpenStudy (anonymous):

yes Jim, and solve for Q

jimthompson5910 (jim_thompson5910):

First, multiply both sides by mtq to get \[\large pmtq=2\] Now divide both sides by pmt to get \[\large q=\frac{2}{pmt}\]

OpenStudy (anonymous):

thankyou for your help in showing the steps, Jim. I appreciate it!